# RiemannFormer: A Framework for Attention in Curved Spaces

### v1

@@ 2 Methodology
$\mathbf{T}_{i}=s^{-\frac{i}{2}}\exp(i\mathbf{X}),$ (7)
@@
$\mathbf{T}_{m}\mathbf{T}_{n}^{-1}$ $=s^{-\frac{m}{2}}\exp(m\mathbf{X})\cdot s^{\frac{n}{2}}\exp(-n\mathbf{X})$ (8)
$=s^{\frac{n-m}{2}}\exp((m-n)\mathbf{X})$
@@
$=\mathbf{q}_{m}^{\top}\left[s^{m}s^{\frac{n-m}{2}}\exp((m-n)\mathbf{X})\right]\mathbf{k}_{n}$
$=\mathbf{q}_{m}^{\top}\left[s^{\frac{m+n}{2}}\exp((m-n)\mathbf{X})\right]\mathbf{k}_{n}$
$=\langle s^{\frac{m}{2}}\exp(-m\mathbf{X})\mathbf{q}_{m},\;s^{\frac{n}{2}}\exp(-n\mathbf{X})\mathbf{k}_{n}\rangle_{\mathbf{I}}$
It should be noted that a skew-symmetric matrix has ${D(D-1)}/{2}$ independent entries and thereby quite a few learnable parameters will be introduced. Similarly, in order to reduce the number of learnable parameters, We can also formally predefine them as block diagonal matrices.
In this work, the notation $\breve{\mathbf{R}}(\theta)$ is used to represent a counterclockwise rotation through an angle $\theta$ in a 2D subspace. We predefine the transformation matrices as follows,
$\mathbf{T}_{m}=s^{-\frac{m}{2}}\begin{bmatrix}\breve{\bf R}(m\theta_{1})&&&\\ &\breve{\bf R}(m\theta_{2})&&\\ &&\ddots&\\ &&&\breve{\bf R}(m\theta_{D/2})\end{bmatrix},$ (10)
and the inverse matrices,
$\mathbf{T}_{m}^{-1}=s^{\frac{m}{2}}\begin{bmatrix}\breve{\bf R}(-m\theta_{1})&&&\\ &\breve{\bf R}(-m\theta_{2})&&\\ &&\ddots&\\ &&&\breve{\bf R}(-m\theta_{D/2})\end{bmatrix}.$ (11)
Then, the product is
$\mathbf{T}_{m}\mathbf{T}_{n}^{-1}=s^{\frac{n-m}{2}}\begin{bmatrix}\breve{\bf R}((m-n)\theta_{1})&&&\\ &\breve{\bf R}((m-n)\theta_{2})&&\\ &&\ddots&\\ &&&\breve{\bf R}((m-n)\theta_{D/2})\end{bmatrix}.$ (12)
By substituting it into Eq. (1), we have
$\langle\mathbf{q}_{m},\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})\rangle_{\mathbf{M}_{m}}=\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})=\mathbf{q}_{m}^{\top}\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}_{n}^{-1}\mathbf{k}_{n}$ (13)
$={\bf q}_{m}^{\top}s^{\frac{n+m}{2}}\begin{bmatrix}\breve{\bf R}((m-n)\theta_{1})&&&\\ &\breve{\bf R}((m-n)\theta_{2})&&\\ &&\ddots&\\ &&&\breve{\bf R}((m-n)\theta_{D/2})\end{bmatrix}\mathbf{k}_{n}.$
It is important to note that the factor $s^{\frac{n+m}{2}}$ disrupts the relativity concerning the positions if $s\neq 1$. Such a mechanism also endows our framework with the capability to balance the capture of both absolute and relative position information. The model autonomously emphasizes the relative information during the learning process by adjusting the parameter $s$. Consequently, this relationship underscores the versatility and broader applicability of our framework. It is also worth noting that, under the aforementioned configuration, the proposed approach reduces to a conventional RoPE when setting $s=1$, thereby indicating that RoPE can be viewed as a special case of our framework.
Furthermore, Eq. (13) can be transformed into the form of a standard inner product,
$\langle s^{\frac{m}{2}}\breve{\bf R}(-m\theta)\mathbf{q}_{m},\;s^{\frac{n}{2}}\breve{\bf R}(-n\theta)\mathbf{k}_{n}\rangle_{\mathbf{I}}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$ (14)
Note that the above formulation represents a simplified version involving only a 2D subspace.
In addition, when considering the transformation matrix $\mathbf{T}_{i}$, we can also use $2\times 2$ reflection matrices as blocks. The notation $\hat{\mathbf{R}}(\theta)$ is used to represent a reflection matrix
$\begin{bmatrix}\cos(2\theta)&\;\;\;\;\!\sin(2\theta)\\ \sin(2\theta)&-\cos(2\theta)\end{bmatrix}.$ (15)
This matrix represents the reflection of an arbitrary two-dimensional vector in a 2D subspace about the axis defined by the vector $\mathbf{d}=(\cos(\theta),\sin(\theta))$. The determinant of a reflection matrix $\hat{\mathbf{R}}(\theta)$ is $-1$, indicating that it reverses orientation. The product of two reflections returned to the original orientation again. Furthermore, this reflection matrix satisfies the property $\hat{\mathbf{R}}(\theta)^{-1}=\hat{\mathbf{R}}(\theta)$, indicating that the matrix is its own inverse. Moreover, the product of rotation matrices and reflection matrices $\widetilde{\bf R}(\theta_{1},\theta_{2})=\hat{\bf R}(\theta_{2})\breve{\bf R}(\theta_{1})$ are still reflection matrices, and the product of two reflection matrices yields a rotation matrix: $\breve{\bf R}(\theta_{1}-\theta_{2})=\hat{\bf R}(\frac{\theta_{1}}{2})\hat{\bf R}(\frac{\theta_{2}}{2})$.
Consequently, beyond rotation matrices, reflection matrices can also be leveraged for the predefinition of the transformation matrix. For example, the rotation matrices in Eq. (10) may be directly substituted as reflection matrices as follows,
$\hat{\mathbf{T}}_{i}=s^{-\frac{i}{2}}\begin{bmatrix}\hat{\mathbf{R}}(i\theta_{1})&&&\\ &\hat{\mathbf{R}}(i\theta_{2})&&\\ &&\ddots&\\ &&&\hat{\mathbf{R}}(i\theta_{D/2})\end{bmatrix}.$ (16)
The subsequent operations follow a similar procedure as previously described and, for the sake of brevity, are omitted here.
We can also predefine the transformations in higher-dimensional subspaces by combining reflections and rotations. Such block matrices as below,
$\hat{\mathbf{R}}(\theta_{1},\theta_{2})=\begin{bmatrix}\breve{\bf R}(\theta_{1})&\mathbf{0}\\[6.79993pt] \mathbf{0}&\hat{\bf R}(\theta_{2})\end{bmatrix}_{4\times 4},$ (17)
and their inverse matrices are
$\hat{\mathbf{R}}(\theta_{1},\theta_{2})^{-1}=\begin{bmatrix}\breve{\bf R}(-\theta_{1})&\mathbf{0}\\[6.79993pt] \mathbf{0}&\hat{\bf R}(\theta_{2})\end{bmatrix}_{4\times 4}.$ (18)
Such combined block matrices may further broaden the space of predefined transformation matrices and allow the model to explore more complex geometric manipulations and representations, thereby gaining deeper insights into the underlying structures and relationships within the data.
2.1 Locality Focusing

### v2

@@ 2.3 A Principled Decomposition of High-Dimensional Rotations
$\mathbf{T}_{i}=\mathbf{M}_{i}^{-\frac{1}{2}}\exp(iX).$ (9)
@@
$\mathbf{T}_{m}\mathbf{T}_{n}^{-1}$ $=\mathbf{M}_{m}^{-\frac{1}{2}}\exp(m{X})\exp(-n{X})\mathbf{M}_{n}^{\frac{1}{2}}$ (10)
$=\mathbf{M}_{m}^{-\frac{1}{2}}\exp((m-n){X})\mathbf{M}_{n}^{\frac{1}{2}},$
@@
$\langle\mathbf{q}_{m},\mathcal{P}_{n\rightarrow m}(\mathbf{k}_{n})\rangle_{\mathbf{M}_{m}}=\mathbf{q}_{m}^{\top}\left[\mathbf{M}_{m}\mathbf{T}_{m}\mathbf{T}_{n}^{-1}\right]\mathbf{k}_{n}=\langle\mathbf{T}_{m}^{-1}\mathbf{q}_{m},\mathbf{T}_{n}^{-1}\mathbf{k}_{n}\rangle_{\mathbf{I}}.$ (11)
It is worth noting that antisymmetric matrices possess D(D1)/2{D(D-1)}/{2} independent elements, thereby introducing a substantial number of learnable parameters. To reduce the number of learnable parameters and mitigate algorithmic complexity, we may alternatively learn subgroups of the orthogonal matrix group $\mathbf{SO}(D)$, such as Maximal Torus $(S^{1})^{k}$, or subspaces of the orthogonal Lie algebra, such as Cartan subalgebras.
We uniformly designate the encoding scheme herein as Parallel Transport of Embeddings (PaTE). When emphasizing a specific subspace, such as SO(4), it is denoted as PaTE-SO(4).
2.3.1 Decomposition into Commuting $\mathbf{SO}(2)$ Subspaces
We can construct a basis for the Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$. This basis consists of $k$ matrices $\{H_{1},H_{2},\dots,H_{k}\}$:
$H_{1}=\begin{bmatrix}J&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{bmatrix},\quad H_{2}=\begin{bmatrix}0&0&\cdots&0\\ 0&J&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\end{bmatrix},\quad\dots\quad H_{k}=\begin{bmatrix}0&0&\cdots&0\\ 0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&J\end{bmatrix}$,
where each $H_{i}$ is a $2k\times 2k$ block matrix, with all blocks being zero matrices except for the $i$-th diagonal block, which is the $2\times 2$ antisymmetric matrix $J$. The matrix $J$ is the generator of infinitesimal rotations in two dimensions:
$J=\begin{bmatrix}0&-1\\ 1&0\end{bmatrix}$.
The Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$ is the linear span of this basis $\{H_{1},H_{2},\dots,H_{k}\}$. Any element $H$ within it can be expressed as:
$H=\sum_{i=1}^{k}\theta_{i}H_{i}=\operatorname{diag}{(\theta_{1}J,\theta_{2}J,\dots,\theta_{k}J)},$
where $\theta_{1},\theta_{2},\dots,\theta_{k}$ are arbitrary real numbers, and $\operatorname{diag}({B}_{1},{B}_{2},\dots,{B}_{n})$ denotes the block diagonal matrix with each ${B}_{i}$ serving as a diagonal block. The matrix $H$ represents simultaneous infinitesimal rotations on $k$ orthogonal planes, with the "angular velocity" on each plane being $\theta_{i}$.
The exponential of $H$ is
$\exp(H)=\exp(\sum_{i=1}^{k}\theta_{i}H_{i})=\operatorname{diag}{(\exp(\theta_{1}J),\exp(\theta_{2}J),\dots,\exp(\theta_{k}J))},$
where a diagonal block $\exp(\theta_{i}{J})$ is a $2\times 2$ rotation matrix,
$\exp(\theta_{i}{J})=\cos(\theta_{i}){I}+\sin(\theta_{i}){J}=\begin{bmatrix}\cos(\theta_{i})&\!\!-\sin(\theta_{i})\\ \sin(\theta_{i})&\;\;\cos(\theta_{i})\end{bmatrix}$.
The interaction between two positions $m$ and $n$ can be represented as
$\exp\left(m\theta{J}\right)\exp\left(n\theta{J}\right)^{-1}$ $=\exp\left((m-n)\theta{J}\right)$
$=\cos\left((m-n)\theta\right){I}+\sin\left((m-n)\theta\right){J}$
$=\begin{bmatrix}\cos((m-n)\theta)&\!\!-\sin((m-n)\theta)\\ \sin((m-n)\theta)&\;\;\cos((m-n)\theta)\end{bmatrix}$.
In fact, for any real antisymmetric matrix $X$, there exists an orthogonal matrix $Q$ such that $X=QHQ^{T}$, where $H$ is an element in the Cartan subalgebra $\mathfrak{h}$ of $\mathfrak{so}(2k)$, and thus $\exp(X)=Q\exp(H)Q^{T}$. Here, the rotation in the 2D subspace corresponds to setting $Q=I$, which simplifies the scenario but also limits its capacity for interactions across spatial dimensions. It is also worth noting that, under the aforementioned configuration, the proposed approach reduces to a conventional RoPE when setting the metric as $\mathbf{M}_{i}=\mathbf{I}$, thereby indicating that RoPE can be viewed as a special case of our framework.
To enhance interactions across more dimensions, one may consider rotational transformations within higher-dimensional subspaces.
2.3.2 $\mathbf{SO}(3)$ Subspaces
We can consider rotational transformations within 3-dimensional subspaces. We assume here that $D$ is a multiple of 3, i.e., $D=3k$.
First, we define the $3\times 3$ unit antisymmetric matrix $K$ corresponding to a unit rotation axis:
$K=\begin{bmatrix}0&-c&b\\ c&0&-a\\ -b&a&0\end{bmatrix}$,
where $a^{2}+b^{2}+c^{2}=1$, i.e., $\mathbf{v}=[a,b,c]^{T}$ is a unit vector representing the rotation axis. For a real number $\theta$, the exponential of the antisymmetric matrix $\theta K$ is:
$\exp(\theta K)=I+\sin(\theta)K+(1-\cos(\theta))K^{2}.$
This is the classic Rodrigues’ rotation formula, which gives a closed-form exponential map from the Lie algebra $\mathfrak{so}(3)$ to the Lie group $\mathbf{SO}(3)$, transforming an antisymmetric matrix encoding the rotation axis and angle into its corresponding rotation matrix.
Incorporating positional information:
$\exp(m\theta K)=I+\sin(m\theta)K+(1-\cos(m\theta))K^{2}.$
Likewise, the interaction between two positions $m$ and $n$ can be represented as
$\exp\left(m\theta K\right)\exp\left(n\theta K\right)^{-1}$ $=\exp\left((m-n)\theta K\right)$
$=I+\sin\left((m-n)\theta\right)K+(1-\cos\left((m-n)\theta)\right)K^{2}.$
2.3.3 The Isoclinic Decomposition of $\mathbf{SO}(4)$
It is well-established that rotations in the special orthogonal group $\mathbf{SO}(4)$ are generated via the exponential map from its Lie algebra, $\mathfrak{so}(4)$. Elements $X\in\mathfrak{so}(4)$ are $4\times 4$ real skew-symmetric matrices of the form:
$X=\begin{bmatrix}0&x_{1}&x_{2}&x_{3}\\ -x_{1}&0&x_{4}&x_{5}\\ -x_{2}&-x_{4}&0&x_{6}\\ -x_{3}&-x_{5}&-x_{6}&0\end{bmatrix}$.
However, for an arbitrary $X$, the direct computation of its matrix exponential, $\exp(X)$, is often complex and lacks a concise closed-form expression analogous to that in 2D, such as $\exp(\theta J)=\cos(\theta)I+\sin(\theta)J$ where $J^{2}=-I$. To devise a more efficient and insightful computational framework, it is imperative to investigate the intrinsic algebraic structure of $\mathfrak{so}(4)$.
The key insight lies in an exceptional property of the $\mathfrak{so}(4)$ Lie algebra: it is the unique semi-simple Lie algebra that decomposes into a direct sum of two ideals. Specifically, $\mathfrak{so}(4)$ is isomorphic to the direct sum of two $\mathfrak{su}(2)$ Lie algebras:
$\mathfrak{so}(4)\cong\mathfrak{su}(2)\oplus\mathfrak{su}(2).$
Given the isomorphism between $\mathfrak{su}(2)$ and the Lie algebra of 3D rotations, $\mathfrak{so}(3)\cong\mathfrak{su}(2)$, the above relation leads to a more intuitive isomorphism over the real field: $\mathfrak{so}(4)\cong\mathfrak{so}(3)\oplus\mathfrak{so}(3)$. This profound structural property implies that the algebra of 4D rotations can be perfectly deconstructed into a combination of two independent, familiar $\mathfrak{so}(3)$ algebras. This foundational decomposition provides the basis for simplifying the exponential map by decomposing $\mathfrak{so}(4)$ into a direct sum of two subalgebras.
To realize this decomposition, we define two special sets of basis vectors in the six-dimensional vector space of $\mathfrak{so}(4)$, denoted $\{A_{1},A_{2},A_{3}\}$ and $\{B_{1},B_{2},B_{3}\}$. These bases respectively span two three-dimensional subalgebras, $\mathfrak{h}_{1}=\text{span}\{A_{1},A_{2},A_{3}\}$ and $\mathfrak{h}_{2}=\text{span}\{B_{1},B_{2},B_{3}\}$. Since their union forms a complete basis for $\mathfrak{so}(4)$, any skew-symmetric matrix $X$ can be uniquely decomposed as $X=A+B$, where $A=\sum a_{i}A_{i}\in\mathfrak{h}_{1}$ and $B=\sum b_{j}B_{j}\in\mathfrak{h}_{2}$. Thus, $\mathfrak{so}(4)$ is the direct sum of its subalgebras $\mathfrak{h}_{1}$ and $\mathfrak{h}_{2}$, denoted $\mathfrak{so}(4)=\mathfrak{h}_{1}\oplus\mathfrak{h}_{2}$.
The principal advantage of this decomposition is the simplification of the exponential map. As any element $A\in\mathfrak{h}_{1}$ commutes with any element $B\in\mathfrak{h}_{2}$, the exponential of their sum factors into a product of exponentials:
$\exp(X)=\exp(A+B)=\exp(A)\exp(B).$
This reduces the computation for a complex 4D rotation to that of two independent, quasi-3D rotations, each admitting an efficient, closed-form solution.
The standard bases for implementing the decomposition $\mathfrak{so}(4)=\mathfrak{h}_{1}\oplus\mathfrak{h}_{2}$ are defined as follows. The basis $\{A_{1},A_{2},A_{3}\}$ for the subalgebra $\mathfrak{h}_{1}$ is:
$A_{1}=\begin{bmatrix}0&0&0&-1\\ 0&0&-1&0\\ 0&1&0&0\\ 1&0&0&0\end{bmatrix},\;A_{2}=\begin{bmatrix}0&0&1&0\\ 0&0&0&-1\\ -1&0&0&0\\ 0&1&0&0\end{bmatrix},\;A_{3}=\begin{bmatrix}0&-1&0&0\\ 1&0&0&0\\ 0&0&0&-1\\ 0&0&1&0\end{bmatrix}$.
Any element in $\mathfrak{h}_{1}$ is a linear combination of these bases. Consider a normalized generator $J_{A}$ defined by a unit vector $(a_{1},a_{2},a_{3})$ such that $a_{1}^{2}+a_{2}^{2}+a_{3}^{2}=1$:
$J_{A}=\sum_{i=1}^{3}a_{i}A_{i}=\begin{bmatrix}0&-a_{3}&a_{2}&-a_{1}\\ a_{3}&0&-a_{1}&-a_{2}\\ -a_{2}&a_{1}&0&-a_{3}\\ a_{1}&a_{2}&a_{3}&0\end{bmatrix}$.
By construction, $J_{A}$ satisfies $J_{A}^{2}=-I$, where $I$ is the identity matrix. Its algebraic behavior is therefore analogous to the imaginary unit $i$. Consequently, its matrix exponential, derived from the Taylor series expansion, assumes a form akin to Euler’s formula:
$\exp(\theta_{1}J_{A})=\cos(\theta_{1}){I}+\sin(\theta_{1})J_{A}.$
Expanding this expression yields an explicit rotation matrix in $\mathbf{SO}(4)$:
$\exp(\theta_{1}J_{A})=\begin{bmatrix}\cos(\theta_{1})&-a_{3}\sin(\theta_{1})&a_{2}\sin(\theta_{1})&-a_{1}\sin(\theta_{1})\\ a_{3}\sin(\theta_{1})&\cos(\theta_{1})&-a_{1}\sin(\theta_{1})&-a_{2}\sin(\theta_{1})\\ -a_{2}\sin(\theta_{1})&a_{1}\sin(\theta_{1})&\cos(\theta_{1})&-a_{3}\sin(\theta_{1})\\ a_{1}\sin(\theta_{1})&a_{2}\sin(\theta_{1})&a_{3}\sin(\theta_{1})&\cos(\theta_{1})\end{bmatrix}$,
where $\theta_{1}$ represents the angle of rotation about an axis defined by the vector $(a_{1},a_{2},a_{3})$.
In a completely analogous manner, we define the second basis $\{B_{1},B_{2},B_{3}\}$, which spans the commuting subalgebra $\mathfrak{h}_{2}$:
$B_{1}=\begin{bmatrix}0&0&0&1\\ 0&0&-1&0\\ 0&1&0&0\\ -1&0&0&0\end{bmatrix},\;B_{2}=\begin{bmatrix}0&0&1&0\\ 0&0&0&1\\ -1&0&0&0\\ 0&-1&0&0\end{bmatrix},\;B_{3}=\begin{bmatrix}0&-1&0&0\\ 1&0&0&0\\ 0&0&0&1\\ 0&0&-1&0\end{bmatrix}$.
We construct a normalized generator $J_{B}$ from a unit vector $(b_{1},b_{2},b_{3})$:
$J_{B}=\sum_{i=1}^{3}b_{i}B_{i}=\begin{bmatrix}0&-b_{3}&b_{2}&b_{1}\\ b_{3}&0&-b_{1}&b_{2}\\ -b_{2}&b_{1}&0&b_{3}\\ -b_{1}&-b_{2}&-b_{3}&0\end{bmatrix}$,
where $b_{1}^{2}+b_{2}^{2}+b_{3}^{2}=1$. The basis $\{B_{i}\}$ shares the same algebraic properties as $\{A_{i}\}$, leading to $J_{B}^{2}=-I$ and the identical form for its exponential map:
$\exp(\theta_{2}J_{B})=\cos(\theta_{2}){I}+\sin(\theta_{2})J_{B},$
where $\theta_{2}$ is another independent angle of rotation.
Rotations generated by $A\in\mathfrak{h}_{1}$, such as $\exp(A)$, are termed left-isoclinic rotations, while those generated by $B\in\mathfrak{h}_{2}$ are right-isoclinic rotations. Any rotation in 4D can be uniquely decomposed into the product of a left-isoclinic and a right-isoclinic rotation, which commute. This gives rise to the exponential property $\exp(A+B)=\exp(A)\exp(B)$ for $A\in\mathfrak{h}_{1},\,B\in\mathfrak{h}_{2}$.
We introduce rotation angles $\theta_{A}$ and $\theta_{B}$ for the two types of rotations. The rotation matrices corresponding to positions $m$ and $n$ are then respectively defined as:
$R(\theta_{A},\theta_{B},m)=\exp\left(m\theta_{A}J_{A}\right)\exp\left(m\theta_{B}J_{B}\right)\quad\text{and}\quad R(\theta_{A},\theta_{B},n)=\exp\left(n\theta_{A}J_{A}\right)\exp\left(n\theta_{B}J_{B}\right).$
The interaction between two positions $m$ and $n$ is captured by the relative transformation:
$R(\theta_{A},\theta_{B},m)R(\theta_{A},\theta_{B},n)^{-1}.$
For a single generator, e.g., $J_{A}$, the relative transformation simplifies to:
$\exp\left(m\theta J_{A}\right)\exp\left(n\theta J_{A}\right)^{-1}=\exp((m-n)\theta J_{A})=\cos((m-n)\theta){I}+\sin((m-n)\theta)J_{A}.$
An identical relation holds for the matrix $J_{B}$. Since the Lie bracket $J_{A},J_{B}]=0$, their exponentials commute, allowing us to write:
$R(\theta_{A},\theta_{B},m)R(\theta_{A},\theta_{B},n)^{-1}=\exp\left((m-n)\theta_{A}J_{A}\right)\exp\left((m-n)\theta_{B}J_{B}\right).$
When the feature dimension $D$ is a multiple of $4k$, we can apply this method to perform rotations in each decoupled 4D subspace. For a position $i$, we define the transformation matrix as:
$\mathbf{T}_{i}=s_{i}^{-1}R(\theta_{A},\theta_{B},i)=s_{i}^{-1}\exp\left(i\theta_{A}J_{A}\right)\exp\left(i\theta_{B}J_{B}\right).$
The interaction between two positions $m$ and $n$ corresponds to the following matrix:
$\mathbf{T}_{m}{\mathbf{T}_{n}}^{\!\!-1}$ $=s_{m}^{-1}R(\theta_{A},\theta_{B},m)s_{n}R(\theta_{A},\theta_{B},n)^{-1}$
$={\frac{s_{n}}{s_{m}}}\exp\left((m-n)\theta_{A}J_{A}\right)\exp\left((m-n)\theta_{B}J_{B}\right).$
2D PaTE-SO(4)
Notably, this approach is applicable not only to 1D token sequences, as in language models, but also to 2D sequences of image patches. We can associate $J_{A}\in\mathfrak{h}_{1}$ with the $x$-axis and $J_{B}\in\mathfrak{h}_{2}$ with the $y$-axis:
$R(\theta_{A},\theta_{B},(x_{1},y_{1}))$ $=\exp\left(x_{1}\theta_{A}J_{A}\right)\exp\left(y_{1}\theta_{B}J_{B}\right),$
$R(\theta_{A},\theta_{B},(x_{2},y_{2}))$ $=\exp\left(x_{2}\theta_{A}J_{A}\right)\exp\left(y_{2}\theta_{B}J_{B}\right).$
For two patches located at $(x_{1},y_{1})$ and $(x_{2},y_{2})$, the relative transformation is:
$\mathbf{T}_{(x_{1},y_{1})}{\mathbf{T}_{(x_{2},y_{2})}}^{\!-1}$ $=s_{x_{1}}^{-1}s_{y_{1}}^{-1}R(\theta_{A},\theta_{B},(x_{1},y_{1}))s_{x_{2}}s_{y_{2}}R(\theta_{A},\theta_{B},(x_{2},y_{2}))^{-1}$
$={\frac{s_{x_{2}}s_{y_{2}}}{s_{x_{1}}s_{y_{1}}}}\exp\left((x_{1}-x_{2})\theta_{A}J_{A}\right)\exp\left((y_{1}-y_{2})\theta_{B}J_{B}\right).$
If we define $\mathbf{M}_{(x_{1},y_{1})}=s^{2}_{x_{1}}s^{2}_{y_{1}}\mathbf{I}$, then the interaction term can be expressed as:
$\mathbf{q}_{\mathbf{m}}^{\top}\mathbf{M}_{\mathbf{m}}\mathbf{T}_{\mathbf{m}}{\mathbf{T}_{\mathbf{n}}}^{\!-1}\mathbf{k}_{\mathbf{n}}$ $=s_{x_{1}}s_{y_{1}}s_{x_{2}}s_{y_{2}}\mathbf{q}_{m}^{\top}\exp\left((x_{1}-x_{2})\theta_{A}J_{A}\right)\exp\left((y_{1}-y_{2})\theta_{B}J_{B}\right)\mathbf{k}_{\mathbf{n}}$
$=\langle\mathbf{T}_{\mathbf{m}}^{-1}\mathbf{q}_{\mathbf{m}},\;{\mathbf{T}_{\mathbf{n}}}^{\!-1}\mathbf{k}_{\mathbf{n}}\rangle,$
where $\mathbf{m}$ and $\mathbf{n}$ denote 2D positions $(x_{1},y_{1})$ and $(x_{2},y_{2})$ for brevity in the expression. The final equation provides a concrete instantiation of our geometric framework for 2D data. The relative transformation is a composition of two orthogonal transformations operating in distinct, commuting subspaces, perfectly encoding the displacement along each spatial axis. This formulation is inherently isotropic, treating the $x$ and $y$ dimensions symmetrically, yet allows for learned anisotropic modulation through the scaling factors $s_{x}$ and $s_{y}$. Ultimately, this method yields a holistic representation of 2D relative positions, standing in contrast to methods that treat spatial dimensions as separate or unstructured features. The unique algebraic structure of $\mathfrak{so}(4)$, specifically its decomposition into two commuting subalgebras, finds a natural correspondence with the geometry of the 2D plane, resulting in a computationally elegant and theoretically robust representation of spatial relationships.
2.4 Inducing Structure on the Trivial Value Bundle
The formalization of value vectors as sections of a trivial bundle, $\mathcal{M}\ \times\mathbb{R}^{D_{v}}$, is not merely a descriptive choice; it is a generative one. It provides a principled foundation for introducing more sophisticated, structured operations beyond simple weighted averaging. We explore three such extensions that leverage the fiber bundle structure to enhance the expressive power of the attention mechanism.
Dynamic Fiber Gating
The standard attention mechanism employs a scalar weight, $a_{m,i}$, which uniformly scales all channels of a value vector $\mathbf{v}_{i}$. This raises a natural question: should the query $\mathbf{q}_{m}$ only determine how much information to retrieve from position $i$, or should it also influence what kind of information is retrieved?
To address this, we introduce a dynamic fiber gating mechanism. We define a learnable gating function, $g$, that is conditioned on the query vector and acts directly on the value fiber $\mathbb{R}^{D_{v}}$. This function maps the query’s tangent space to a gating vector in the value space:
$g:T_{\mathbf{p}_{m}}\mathcal{M}\rightarrow\mathbb{R}^{D_{v}}.$
The aggregation of values is then modulated by this gate, allowing for content-addressable channel selection:
$\mathbf{o}_{m}=\sum_{i}a_{m,i}\left(g(\mathbf{q}_{m})\odot\mathbf{v}_{i}\right),$
where $\odot$ denotes the element-wise product. This mechanism empowers the model to dynamically select and emphasize the most relevant feature channels from a value vector, conditioned on the specific query.
Local Symmetry via Group Action
The trivial bundle construction assumes that value vectors can be directly aggregated. However, for data with inherent geometric properties, it may be beneficial to apply a position-dependent transformation to "align" features before aggregation.
This can be formalized by introducing a group action on the value fiber. Let $G$ be a Lie group acting on $\mathbb{R}^{D_{v}}$, such as the special orthogonal group $\mathbf{SO}(D_{v})$. We can then learn a mapping, $h$, from the base manifold to the group itself:
$h:\mathcal{M}\rightarrow G.$
The value aggregation is then modified to incorporate these local symmetry transformations:
$\mathbf{o}_{m}=\sum_{i}a_{m,i}\left(h(\mathbf{p}_{i})\cdot\mathbf{v}_{i}\right),$
where $h(\mathbf{p}_{i})\cdot\mathbf{v}_{i}$ represents the action of the group element $h(\mathbf{p}_{i})$ (e.g., a rotation matrix) on the vector $\mathbf{v}_{i}$. This introduces a powerful inductive bias, often referred to as a local gauge symmetry, allowing the model to learn position-aware feature adaptations.
Connections on the Value Bundle
Our initial assumption of a trivial bundle implies that the "rules" for aggregating values are globally uniform and position-agnostic. We can challenge this assumption by introducing a new, learnable connection on the value bundle.
A connection defines a notion of parallel transport specific to the bundle it is on. Let us denote this new value connection as $\mathcal{C}$, with its corresponding parallel transport operator $\mathcal{C}_{i\rightarrow m}$. Crucially, this connection is independent of the Riemannian connection on the tangent bundle used for $Q/K$ interactions. The final aggregation step then becomes:
$\mathbf{o}_{m}=\sum_{i}a_{m,i}\,\mathcal{C}_{i\rightarrow m}(\mathbf{v}_{i}).$
This establishes a dual-geometry framework:
1) A Riemannian geometry on the tangent bundle governs the $Q/K$ interactions, determining where to attend.
2) A gauge-like geometry, defined by the learned value connection $\mathcal{C}$, governs the $V$ aggregation, determining how to integrate the attended information.
This provides maximal flexibility, allowing the model to learn, for instance, that information from distant tokens should be transformed (e.g., blurred or down-weighted) differently than information from nearby tokens during the final value aggregation.
A Concrete Realization: The Locality Focusing Mechanism
While the concepts of dynamic gating, group actions, and fully learned connections represent powerful avenues for future research, we now introduce a practical and highly effective mechanism that realizes the core principle of a structured value bundle with minimal overhead. This approach can be understood as a specific instantiation of a predefined connection on the value bundle. Instead of a complex learnable operator, this connection is simplified to a position-dependent scalar modulation, designed explicitly to enforce a local inductive bias. We term this mechanism Locality Focusing.